OpenStudy (anonymous):

Question edited : I only don't know how to integrate x/(x^2+x+1) Please help with this

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=x%2F%28x^3-1%29

OpenStudy (anonymous):

\[ \frac x{x^3-1}=\frac { A x + B}{x^2+ x+1}+ \frac C {x-1} \]

OpenStudy (anonymous):

yeah but it does't work , cant even integrate a nimber like that the first term Ax+B

OpenStudy (anonymous):

Multiply the two sides by x-1, you get \[ \frac {x(x-1)}{x^3-1}=\frac {( A x + B)(x-1)}{x^2+ x+1}+ \frac {C(x-1)} {x-1}\\ \frac {x}{x^2+ x+1}=\frac {( A x + B)(x-1)}{x^2+ x+1}+ C \] Put x=1 you get\[ \frac 1 3= 0 + C\] We have C =1/3

OpenStudy (anonymous):

yeah c=1/3 b=-1/3 and A=C

OpenStudy (anonymous):

Now multiply the two sides by x and let x goes to infinity \[ \frac {x^2}{x^3-1}=\frac { A x^2 + Bx}{x^2+ x+1}+ \frac x {3(x-1) }\\ 0=A + \frac 13\\ A=-\frac 13 \]

OpenStudy (anonymous):

Now \[ \frac x{x^3-1}=\frac { A x + B}{x^2+ x+1}+ \frac1 {3(x-1)} \\ x=0 \\ 0 =\frac { + B}{+1}+ \frac1 {3(0-1)}\\ B=\frac 13 \]

OpenStudy (ranga):

I think eldi has already done the partial fractions as shown on the third line of the problem.

OpenStudy (anonymous):

I did it by what we call the cover-up method \[ \frac{x}{x^3-1}=\frac{1-x}{3 \left(x^2+x+1\right)}+\frac{1}{3 (x-1)} \]

OpenStudy (anonymous):

this is the same answer i got

OpenStudy (anonymous):

but i can't integratate this fraction still

OpenStudy (ranga):

\[\frac{1-x}{3 \left(x^2+x+1\right)} = \frac{ 1 }{ 3 } * [ \frac{3}{2 \left(x^2+x+1\right)} - \frac{2x+1}{2 \left(x^2+x+1\right)} ]\]

OpenStudy (anonymous):

i am not sure which is right but the part (x^2 +x+1 ) i still can not integrate that ???

hartnn (hartnn):

a general rule, \(\\ \text{ To integrate }\quad \huge \frac{ax+b}{\sqrt{px^2+qx+r}} or \frac{ax+b}{px^2+qx+r} \\ \large express \quad ax+b=A\frac{d}{dx}(px^2+qx+r)+B \\ \text{then separate D.} \)

hartnn (hartnn):

**separate the Denominator

hartnn (hartnn):

for 1/(x^2+x+1) try to complete the square in the denominator

OpenStudy (anonymous):

i am trying completing the square ,

hartnn (hartnn):

you separated (1-x)/(denom) = 1/ denom - x/ denom ?? don't do that, do what ranga proposed. did you get what he did ?

hartnn (hartnn):

\(\frac{1-x}{3 \left(x^2+x+1\right)} = \frac{ 1 }{ 3 } * [ \frac{3}{2 \left(x^2+x+1\right)} - \frac{2x+1}{2 \left(x^2+x+1\right)} ]\)

OpenStudy (anonymous):

I changed the question on top, that is the only thing i don't get from the problem

hartnn (hartnn):

ok then. since d/dx (x^2+x+1) is 2x+1 we write "x" in numerator as (1/2) (2x) = (1/2) (2x+1 - 1) = (1/2) (2x+1) - 1/2 then separate the denominators.

OpenStudy (anonymous):

Thank you very much, you see I am have a test on two days and I have been studying all day couldn't get around this one than you again

hartnn (hartnn):

oh, so you could solve it entirely?! ask if you atill have any doubt in any step :)

OpenStudy (anonymous):

no no it's ok i get it now ,,but i have never seen that type of manipulation before the one you wrote

hartnn (hartnn):

ohh, with such types, linear / quadratic or linear/ sqrt (quadratic) such manipulation is quite common :) good luck for your test!

OpenStudy (anonymous):

Here is a similar solved problem. There is an integral sign missing in the line before last. This problem was generated by http://www.saab.org/calculus.cgi and selecting Techniques of Integrations.